Home‎ > ‎Crew 141‎ > ‎

502-AstronomyReport

Astronomy Report
05/02/2014
Pete Morgan-Dimmick

Sky Conditions:
Clear
Wind Conditions:
Calm
Observation Start Time:
21:00
Observation End Time:
22:30
Summary:
Unable to get the telescope to align properly, gave up after an hour or two. Will try again tonight.
Objects Viewed:
none
Problems Encountered:
Telescope alignment issues. Email from last night has all the details. Will try again tonight.

Pete Morgan-Dimmick
MDRS Crew 141
XO / Acting Commander